SYMPATHETICAL LASER COOLING OF MOLECULAR IONS TO THE μK REGIME

\underline{C. RICARDO VITERI} and KENNETH R. BROWN, School of Chemistry and Biochemistry and Computational Science and Engineering Division, Georgia Institute of Technology, Atlanta, Georgia 30332, USA.

The complexity of molecular spectra prevents direct laser cooling of most molecules. Molecular ions can be indirectly laser-cooled by a Columbic interaction with a neighboring atomic ion. This sympathetic cooling method has been used to lower the temperature of molecular ions ranging from CaO^{+} to C_{60}^{+} to less than 100 mK by Doppler cooling the atomic ions^{a}. Starting with two ions of Ca^{+} in the trap, we introduce oxygen gas until one CaO^{+} is produced. The motion of a CaO^{+} molecular ion and a Ca^{+} atomic ion are coupled by the Coulomb interaction in the same trap. The frequencies of the normal modes of the atom-molecule crystal are first measured by observing the fluorescence quenching of the Ca^{+} atoms. The normal modes can be cooled to the motional ground state by addressing the sidebands of the Ca^{+} quadrupole transition. The result is an atom-molecule crystal with a translational temperature in the μK regime.

aM. Drewsen, A. Mortensen, R. Martinussen, P. Staanum and J. L. Sørensen Phys. Rev. Lett. \underline{93}, 243201 (2004). V. L. Ryjkov, XZ. Zhao and H. A. Schuessler Phys. Rev. A \underline{74}, 023401 (2006). A. Ostendorf, C. B. Zhang, M. A. Wilson, D. Offenberg, B. Roth, and S. Schiller Phys. Rev. Lett. \underline{97}, 243005 (2006).